OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

-three fourthsx + y = 10 3x - 4y = -40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

equation editor really helps for these lazy is no excuse

OpenStudy (anonymous):

\[y-4=\frac{3}{4}(x+8)\] is not so hard to write solve this for \(y\) in two steps

OpenStudy (anonymous):

multiply out using the distributive law on the left \[y-5=\frac{3}{4}x+6\] then add \(5\) to get \[y=\frac{3}{4}x+11\]

OpenStudy (anonymous):

Omg thank you

OpenStudy (anonymous):

I understand it now